QueenMidnight16:

Solve x3 = 1 over 8. 1 over 2 ±1 over 2 1 over 4 ±1 over 4

SmokeyBrown:

\[x ^{3} = \frac{ 1 }{ 8 }\] Yes? In this case, "x" is a number that, when multiplied by itself 3 times gives us 1/8; you could also say x*x*x = 1/8 We can try each of the options in turn to see if they work. I'll start from the bottom so as not to give away the answer too quickly: \[\frac{1}{4} * \frac{1}{4} * \frac{1}{4} = \frac{1}{64}\] So neither the fourth nor third options would be correct Now, \[\frac{1}{2} * \frac{1}{2} * \frac{1}{2} = \frac{1}{8}\] but \[-\frac{1}{2} * -\frac{1}{2} * -\frac{1}{2} = -\frac{1}{8}\] So this would not be correct Only POSITIVE 1/2 works in this case, NOT NEGATIVE 1/2, since you get a different result for -1/2 I trust that you can see which answer you should choose, yes?

QueenMidnight16:

oooh i get it now. but i dont understand this ±

SmokeyBrown:

That symbol, ± is the same as saying + AND - So, ±1/2, for example, would include positive 1/2 and negative 1/2 in the solution; ±1/4 would be the same as saying positive 1/4 and negative 1/4, etc. You use ± when you want to say both the positive and the negative values of the number

QueenMidnight16:

ok

QueenMidnight16:

thank you